OpenStudy (anonymous):

what is cos 2pi/3? a. -3sqrt/2 b. sqrt 3/2 c. -1/2 d. 1/2 ( i think its b or c)

OpenStudy (kc_kennylau):

Convert it into degrees if that helps; or draw a unit circle if that helps

OpenStudy (kc_kennylau):



OpenStudy (anonymous):

120'

OpenStudy (kc_kennylau):

What's its supplementary angle?

OpenStudy (kc_kennylau):

Formula: \(\cos(180^\circ-\theta)=-\cos\theta\)

OpenStudy (anonymous):

60?

OpenStudy (kc_kennylau):

Or you can use \(\cos(90^\circ+\theta)=-\sin\theta\) which is more standard

OpenStudy (kc_kennylau):

Choose one of the formulas that I gave you

OpenStudy (anonymous):

the first one.

OpenStudy (kc_kennylau):

Then use it

OpenStudy (anonymous):

but what do i use in place of the theta?

OpenStudy (kc_kennylau):

60`

OpenStudy (anonymous):

in both of them?

OpenStudy (kc_kennylau):

Both of what? You've chosen a formula already

OpenStudy (anonymous):

yes. but do i put 60 in both of the thetas?

OpenStudy (kc_kennylau):

\[\cos120^\circ=\cos(180^\circ-60^\circ)=\cdots\]

OpenStudy (kc_kennylau):

P.S. Schools prefer the second formula

OpenStudy (kc_kennylau):

So I'll use the first formula and you'll try to use the second formula. Using the first formula: \(\cos120^\circ=\cos(180^\circ-60^\circ)=-\cos60^\circ=-0.5\)

OpenStudy (anonymous):

-1/2?

OpenStudy (kc_kennylau):

Yes

OpenStudy (kc_kennylau):

Standard method: \(\displaystyle\cos\frac{2\pi}3=\cos\left(\frac\pi2+\frac\pi6\right)=-\sin\left(\frac\pi6\right)=-\frac12\)